OpenStudy (anonymous):

Evaluate: -6 - (-4) ? Evaluate: -14 + 8 ?

OpenStudy (merchandize):

-6+4 =-2 -14+8 =-6

OpenStudy (anonymous):

Do you know what happens when you have double minussigns?

OpenStudy (anonymous):

-2 for number one and =6 for number two exatly ^^

OpenStudy (anonymous):

-6**

OpenStudy (anonymous):

thanksss guyss (: & frx noppp im lost in math

OpenStudy (anonymous):

That's a good thing to remeber. Then you get -(-) "minus minus" it changes into plus

OpenStudy (hba):

Oh My God So Diff Questions ?